Recreational vehicles use propane for cooking and heating. The combustion of propane and oxygen produces carbon dioxide and water.

What STP volume of carbon dioxide gas is produced from the combustion of 14.0 g of propane according to the combustion reaction?
Answer #1

Mass of propane combusted = 14.0 g

Molar mass of propane = 44.1 g/mol

Moles of propane combusted = 14.0 g/44.1 g/mol = 0.317 mol

C3H8 + 5O2 3CO2 + 4H2O

Now, from the above balanced equation, it is clear that 1 mol of propane on combustion gives 3 mol of carbon dioxide.

Therefore, 0.317 mol of propane on combustion gives (3 mol x 0.317 mol)/1 mol = 0.951 mol of carbon dioxide.

At STP, 1 mol carbon dioxide occupies 22.4 L

Therefore, 0.951 mol of carbon dioxide occupies = (0.951 mol x 22.4 L)/1 mol = 21.3 L

Hence, volume of carbon dioxide gas is produced = 21.3 L